OpenStudy (selena2345):

Where do the lines modeled by the equations y=-7/8x+1/4 and y=-3/4x+1/2 intersect? A:(-4,2) B:(-2,2) C:(2,-2) D:(4,-2)

OpenStudy (bahrom7893):

what they are asking you is: What are x and y when: -7/8x+1/4 = -3/4x+1/2

OpenStudy (selena2345):

yes

OpenStudy (bahrom7893):

solve for x, then plug whatever value you get back into either of those equations to get y. That will be your answer.

OpenStudy (selena2345):

can u show me please ?

OpenStudy (bahrom7893):

Multiply everything by 8: -7x + 2 = -6x + 4 Move the terms with x to one and the rest to the other side: -7x + 6x = 4 - 2 Simplify: -x = 2 x = -2

OpenStudy (selena2345):

can give me an example of that ?

OpenStudy (bahrom7893):

y=-7/8x+1/4 = -7*-2/8 + 1/4 = 14/8 + 1/4 = 7/4 + 1/4 = 8/4 = 2

OpenStudy (bahrom7893):

(x, y) = (-2, 2)

OpenStudy (selena2345):

thank you @bahrom7893